OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = -x6 - x5- x4 - 4x3 - 12x2 + 12 ? (1 points) Answer Positive: 4, 2, or 0; Negative: 2 or 0; Complex: 6, 4, 2, or 0 Positive: 3 or 1; Negative: 3 or 1; Complex: 4, 2, or 0 Positive: 3 or 1; Negative: 2 or 0; Complex: 4, 2, or 0 Positive: 1; Negative: 5, 3, or 1; Complex: 4, 2, or 0

OpenStudy (anonymous):

My answer is B

OpenStudy (amoodarya):

x1 = 0.8371 x2 = -1.1577 x3 = -1.336 + 1.1789i x4 = -1.336 - 1.1789i x5 = 0.9963 - 1.7053i x6 = 0.9963 + 1.7053i

OpenStudy (amoodarya):

d is right answer but also ithink that question has a logic error

OpenStudy (anonymous):

Slightly confused

OpenStudy (anonymous):

oh, i see....thank you kind sir!!!!

OpenStudy (amoodarya):

what method you have to apply ? for example bairstow method , bisection , ...